OpenStudy (anonymous):

The number of visitors to a park is expected to follow the function v(x) = 7(x − 1), where x is the number of days since opening. On the first day, there will be a ceremony with 39 people in attendance. What is the function that shows total visitors, including the ceremony? @math&ing001

OpenStudy (anonymous):

Help again please @paki @math&ing001 @midhun.madhu1987

OpenStudy (math&ing001):

On the first day v(1)=0 so the total number of visitors will be represented by f(x) = v(x) +39 = 7(x-1) +39

OpenStudy (anonymous):

Can you help me with another please